a. Following information is available in respect of a company named, as Beta Ltd
Particulars Amount
Profit before taxes Rs 50 lakhs
Dividend declared per share Rs2
Price of the share prevalent on stock
exchange Rs 200
Applicable tax rate 35%
Share capital of company 5 lakh shares
of Rs 10 each
Calculate-
•Earnings per share
•Dividend Pay-out Ratio
•Price earnings ratio
1
Expert's answer
2020-01-14T11:01:19-0500

Earnings per share 


Earnings per share or basic earnings per share is calculated by subtracting preferred dividends from net income and dividing by the weighted average common shares outstanding. The earnings per share formula looks like this.


Earning per share =(Net income -prefered dividends)/Weighted Average Common shares outstanding.

=65% of profit before taxses=Net income=35

prefered dividens = 2

Averarge = 10


Therefore ; Earning per share= (35-2)/10

EPS= 33/10

=3.3


Dividend Pay-out Ratio


The dividend payout ratio can be calculated as the yearly dividend per share divided by the earnings per share, or equivalently, the dividends divided by net income (as shown below)

"\u200b\t \nDividend Payout Ratio= \nNet Income\/\nDividends Paid\n\u200b\t \n\u200b"


=65% of profit before taxses=Net income=35

=Dividend declared per share Rs2 x5=10

therefore ;dividend payout ratio= 35/10x100=35%

=35%


Price earnings ratio



The P/E ratio is calculated by dividing a company's current stock price by its earnings per share (EPS). If you don't know the EPS, you can calculate it by subtracting a company's preferred dividends paid from its net income, and then dividing the result by the number of shares outstanding.


Price of the share prevalent on stock exchange Rs 200 

Earning per share =(Net income -prefered dividends)/Weighted Average Common shares outstanding.

=65% of profit before taxses=Net income=35

prefered dividens = 2

Averarge = 10


Therefore ; Earning per share= (35-2)/10

EPS= 33/10

Price earnings ratio

=3.3


200/35=5.7143